A joint survey by a magazine and a website found that 57% of workers say that if they could do it all over again, they would choose a different career. The survey also found that 33% of workers say they plan to retire early and 67% say they plan to wait and retire at age 65 or older. Assume that the following joint probability table applies.

Retire Early
Yes No
Career Same 0.21 0.22 0.43
Different 0.12 0.45 0.57
0.33 0.67

(a)

What is the probability a worker would select the same career?

(b)

What is the probability a worker who would select the same career plans to retire early? (Round your answer to four decimal places.)

(c)

What is the probability a worker who would select a different career plans to retire early? (Round your answer to four decimal places.)

(d)

What do the conditional probabilities in parts (b) and (c) suggest about the reasons workers say they would select the same career?

The probability the worker plans to retire early is  ---Select--- greater smaller for the worker who would choose the same career. This suggests that one of the reasons the worker would select the same career is that there is a  ---Select--- greater smaller probability of being able to retire early. On the other hand, the worker who would choose a different career may have a  ---Select--- greater smaller probability of being able to retire early.
Answer #1

Solution:-

(a)

The probability a worker would select the same career is=0.43

(b)

P(Retire Early |Same Career) = P(Retire Early and Same Career) / P(Same Career) = 0.21 /0.43 = 0.4884

(c)

The probability a worker who would select a different career plans to retire early is

P(Retire early |Different career) = P(Retire early and different career) / P(Different career) = 0.12 /0.57 = 0.2105

I have solved all the parts please fill the part d. as drop down is not given, or provide the drop down i will help you. also if there is any confusion to understand any part please comment.